OpenStudy (anonymous):

It took Smedly 5 and 3/4 hours to climb to the top of the mountain it took 3 and 1/4 hours to climb down if he spent 1 and1/2 hours at the top how long did the climb take? And explain please

OpenStudy (anonymous):

(5 + 3/4 ) + (3 + 1/4 ) = 9 hr (9 hr ) + (1 + 1/2 hr) = 10 and 1/2 hr

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

\[Time_{total} = Time_{up} + Time_{resting} + Time_{down}\] \[=5 + 3/4 + 3 + 1/4 + 1 + 1/2\] Since we are free to move around terms in a sum... \[=(5+3) + (3/4 + 1/4) + 1/2\] \[= 8 + 4/4 + 1/2\] \[= 8 + 1 + 1/2\] \[= 9 + 1/2\] Therefore the trip took 9 and 1/2 hours.

OpenStudy (anonymous):

Ack. I forgot a 1. Should be 10 and a half.

OpenStudy (anonymous):

5 + 3 + 3/4 +1/2 = 9hr

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

Also I had the times in the wrong order. But did you understand what I meant?

OpenStudy (anonymous):

Yes thanks